# Measuring Dwarf Galaxy Intrinsic Abundance Scatter with Mid-resolution Spectroscopic Surveys: Calibrating APOGEE Abundance Errors

### v1

@@ 3.2.1 Reported APOGEE Uncertainties
[…]
@@ 3.2.2 High-Fidelity G23 Data Sample
[…]
Table 1: Number stars in each calibration bin.
-2.1 < [Fe/H] < -1.55 -1.55 < [Fe/H] < -1
25 < S/N < 125 21 84
125 < S/N < 225 35 134
225 < S/N < 325 27 109
325 < S/N < 425 17 52
425 < S/N 36 113
[…]
Figure 6: Added error terms for all elemental abundances, binned by SNR and [Fe/H]. Within each elemental abundance, added error terms generally decrease with increase SNR and [Fe/H].
G23 find intrinsic scatters ranging from $0.04-0.08$ dex in [X/Fe] for element abundances in our set (Mg, Si, Ti, Ni, Mn). This is greater than the photon-noise uncertainty by a factor of 2 for most elements. Overall, the errors that G23 calculate for their data range from an average of 0.048 dex on [Si/Fe] to 0.068 dex on [Ti/Fe]. Because the uncertainties are significantly smaller compared to the scatter, small inaccuracies in the abundance uncertainty will have a small impact on the intrinsic scatter. Therefore, the method provides robust upper limits on the intrinsic scatter of the element abundances for this population.
To be consistent with the calculated APOGEE intrinsic scatters, we use Equation 2 to calculate the intrinsic scatter of the G23 data for the 86 stars in their sample, showing the result in Figure 5. We report a range of intrinsic scatters in [X/Fe] of $0.04-0.08$ dex for elements that overlap with APOGEE (Mg, Si, Ti, Ni, Mn). The calculated intrinsic scatters for G23 are within (Mg, Si, Ti) or slightly lower (Ni, Mn) than the reported intrinsic scatters relative to the two- and three-parameter models reported in Table 4 of G23.
@@ 3.2.3 Calibration of APOGEE Uncertainties
Due to the robustness of intrinsic scatter from the G23 data, we take these measurements as the "ground truth" intrinsic scatter of halo field stars and calibrate the uncertainties on the APOGEE data such that the calculated intrinsic scatter of the APOGEE halo data matches that of G23. In practice, we calculate an additional term, $a$, every abundance and add it in quadrature to the reported uncertainty on abundance measurements in APOGEE.
@@
Using equation 3 we fix the intrinsic scatter, $s$, to the calculated value from the G23 data and maximize the likelihood over $a$. We bin over the APOGEE halo data in [Fe/H] and SNR as we expect that the total uncertainties are primarily dependent on these quantities. Due to the small number of high SNR stars at low metallicity, we divide our [Fe/H] range into two bins from $-2.1$ dex to $-1.55$ dex and $-1.55$ dex to $-1$ dex. We then bin the SNR in ranges of 100, from an SNR of 25 up to 425. The final bin contains all stars with SNR higher than 425 as there are very few stars in this range and SNRs beyond this range are thought to be unphysically high due to persistence effects (Holtzman et al., 2018), though the data does still have very high SNR. We note that finer binning would result in more precise added error terms on the data, but that we are limited by the APOGEE halo sample size. The number of stars in each bin is listed in Table 1.
@@
25 < S/N < 125 125 < S/N < 225 225 < S/N < 325 325 < S/N < 425 425 < S/N
[Al/Fe], [O/Fe] -2.1 < [Fe/H] < -1.55 0.17 0.14 0.16 0.09 0.08
-1.55 < [Fe/H] < -1 0.22 0.10 0.03 0.03 0.04
[Mg/Fe] -2.1 < [Fe/H] < -1.55 0.33 0.09 0.00 0.00 0.00
-1.55 < [Fe/H] < -1 0.24 0.10 0.00 0.00 0.00
[Si/Fe] -2.1 < [Fe/H] < -1.55 0.27 0.18 0.19 0.00 0.00
-1.55 < [Fe/H] < -1 0.25 0.09 0.00 0.00 0.06
[Ti/Fe] -2.1 < [Fe/H] < -1.55 0.12 0.11 0.29 0.27 0.19
-1.55 < [Fe/H] < -1 0.20 0.15 0.05 0.00 0.06
[Ni/Fe] -2.1 < [Fe/H] < -1.55 0.01 0.21 0.14 0.10 0.07
-1.55 < [Fe/H] < -1 0.17 0.07 0.00 0.00 0.00
[Mn/Fe] -2.1 < [Fe/H] < -1.55 0.13 0.11 0.17 0.05 0.13
-1.55 < [Fe/H] < -1 0.22 0.09 0.12 0.12 0.08
Figure 6 shows the added error terms for all abundances explored with the APOGEE data. There is a clear trend of decreasing added error terms with higher SNR and [Fe/H]. This is expected as higher SNR and higher metallicity data have stronger metal lines. Furthermore, we note that the size of the added error terms at low SNR are on the order of the uncalibrated intrinsic scatter. This implies the intrinsic scatters are very sensitive to the estimated uncertainty which is a limitation of the data quality, and that the intrinsic scatters reported in this study are upper bounds. As we construct the G23 halo and APOGEE samples to be drawn from nearly the same underlying population, this is a robust data-driven method to obtain accurate uncertainties that we can propagate for our analysis of the dSph systems and provides a benchmark for comparison for further studies.
@@ 4.1 Milky Way Stellar Halo
To validate our calibrated uncertainty estimates, we first compare the calibrated intrinsic scatter of the APOGEE halo sample to the G23 halo sample as shown in Figure 8. Calibrated intrinsic scatters for all abundances decrease by at least 0.02 dex compared to the uncalibrated intrinsic scatters. For those abundances overlapping with the G23 data, intrinsic scatters for the calibrated APOGEE data are approximately the same as the intrinsic scatters derived from the G23 data, as expected. Specifically, we report intrinsic scatters of 0.07 dex for [Mg/Fe], 0.06 dex for [Si/Fe], 0.06 dex for [Ti/Fe], and 0.05 dex for [Ni/Fe], which are all within 0.015 dex of G23. We report an intrinsic scatter of 0.11 for [Mn/Fe], notably 0.03 dex higher than the G23 value, but this is perhaps due to the added error terms for [Mn/Fe] having greater variation particularly in the high-metallicity bin. This suggests that the data may need to be more finely binned to get an accurate estimate of the added error for each star, which would require a larger sample of low-metallicity stars. In general, the differences compared to G23 are likely due to the coarse binning for the added error term. In addition, since the maximization of the joint probability function accounts for the individual errors on stars, and the added error terms are not calibrated for individual stars; this likely introduces some small differences between the intrinsic scatter in the G23 data and the calibrated APOGEE data. Nevertheless, these differences are slight and the comparison in Figure 8 validates our procedure.
@@ 4.2 Milky Way dSphs
With the robustness of our calibration confirmed, we now turn to an analysis of the dSphs. Figure 7 shows the comparison between the calibrated and uncalibrated average dSph intrinsic scatter, calibrated intrinsic scatters for individual dSphs in APOGEE, and the average intrinsic scatters from the K18 sample. The calibrated average intrinsic scatter of dSphs is reduced by up to 0.11 dex overt the uncalibrated data, making the average scatter of dSphs in the APOGEE data comparable to the average scatter of dSphs in K18.
The data also shows that generally, the dSphs have lower intrinsic scatter and dispersion in the light and $\alpha$-elements (average $0.04\pm 0.02$ dex) than in the Fe-peak elements (average $0.11\pm 0.05$ dex). This variation may arise from the difference in time-scales for core-collapse versus Type Ia SNe. More common core-collapse SNe occur relatively frequently and on short-timescales which could result in greater mixing over time and lower impact from any single event, leading to lower intrinsic scatter. Conversely, less frequent longer-timescale Type Ia SNe could result in less mixing before star formation episodes, thereby increasing intrinsic scatter.
@@ 4.3 Milky Way Globular Clusters
Figure 9 shows the average intrinsic scatters of 1604 stars across 19 GCs (median 79 stars per GC) that were targeted in APOGEE DR17 (see the Appendix for a list of GCs). We separately analyze Omega Centauri and M54, as they are suspected cores of disrupted dwarf galaxies (Neumayer et al., 2020). The GCs have a median SNR of 157, and [Fe/H] of -1.1. The average intrinsic scatter is $0.21$ dex for [Al/Fe] and $0.08$ for [O/Fe]; higher than the $\alpha$-elements which have an average scatter of $0.04$ dex and the Fe-peak elements which have an average scatter of 0.05 dex. This is consistent with the multiple stellar populations mentioned previously.
@@ 5.1 Summary
In this work we introduce a methodology for calibrating abundance uncertainties derived from mid-resolution surveys by leveraging smaller, high-fidelity, high-resolution data sets. Data-driven calibration or label-transfer efforts have previously been employed to garner greater value from low SNR data (e.g. Casey et al., 2017; Ho et al., 2017; Rice & Brewer, 2020; Xiang et al., 2021, 2022). However, these typically rely on stars in common between the surveys. Our approach does not rely on this, but instead utilizes population statistics. We apply our method to calibrate the measurement errors and calculate the intrinsic abundance scatter of metal-poor stars for elements Al, O, Mg, Si, Ti, Ni, and Mn in dSphs, GCs, and the Milky Way stellar halo. Our calculation relies on the calibration of the reported APOGEE abundance uncertainties and is specifically for APOGEE DR17 metal-poor red giant stars ($-2.1\leq$[Fe/H]$\leq-1$) and in the temperature range 4000KT $4000\,\rm{K}\,\leq T\leq\,5000\,\rm{K}$.
@@
Making measurements of the intrinsic abundance scatter of dSphs, GCs, and the halo with the APOGEE data, particularly for metal-poor stars, is important for providing constraints for simulations and linking to physical drivers. Doing this requires using data at lower metallicities and SNR than the majority of APOGEE spectra, in a regime where there is little previous work. As the magnitude of the uncertainties impact the measurement of intrinsic scatter, it is vital to test the accuracy of the APOGEE reported uncertainty. In doing so, we found them to be underestimated and using the method introduced in this paper, we derive the complete uncertainties on the APOGEE data in this regime.
@@
To do this, we use a high fidelity reference set of halo stars (G23) to provide ground truth information about the intrinsic scatter of element abundances in the stellar halo. We construct a comparable sample of field halo stars from the APOGEE data using the same distance and metallicity criteria as G23. Our primary assumption in this work is that the underlying population of stars in the reference set (G23) is comparable to the halo population in the calibration set we use from APOGEE. Comparing the intrinsic scatters between the two datasets, we find that the intrinsic scatters of the uncalibrated APOGEE halo data are substantially higher than G23, indicating the APOGEE uncertainties are underestimated. This is unsurprising as the reference objects used in APOGEE for providing empirical uncertainty estimates are sparse in the metal-poor and low SNR regime (SNR $\sim$ 50) (Jönsson et al., 2020). We note that the G23 uncertainties themselves are very small so their intrinsic scatters, which we take as ground truth, are minimally impacted by any inaccuracy in these. However, strictly, any underestimation in the G23 uncertainties would result in lower intrinsic scatter, and so these data set an upper limit on the ground truth intrinsic scatter and therefore a lower limit on the complete uncertainties for APOGEE.

### v2

@@ 3.2.1 Reported APOGEE Uncertainties
Uncertainties on the derived parameters are estimated from the inverse of the curvature matrix which depends on the flux error and partial derivatives of the synthetic spectra with the different parameters estimated. These uncertainties are likely underestimated. Indeed, estimates of abundance scatter in star clusters, yield substantially larger errors than the internal FERRE uncertainty estimates (Holtzman et al., 2015). These uncertainties were adjusted in recent data releases using repeat observations, but abundance measurements may still have systematic errors for metal-poor stars as the data is sparse in this regime and at low S/N (Jönsson et al., 2020). Repeat observations provide an empirical estimate of uncertainty as a function of Teff, [M/H] and S/N (Holtzman et al., in prep). These empirically derived uncertainties are reported for each abundance in the data release table which we used to calculate the intrinsic scatters in Figures 4 and 5.
@@
Table 1: Median number of stars in each calibration bin.
-2.1 < [Fe/H] < -1.55 -1.55 < [Fe/H] < -1
25 < S/N < 125 9 46
125 < S/N < 225 9 51
225 < S/N < 325 11 53
325 < S/N < 425 8 24
425 < S/N 19 50
@@ 3.2.2 High-Fidelity G23 Data Sample
The primary goal of the survey was to measure the intrinsic abundance scatter. To do so, they selected a stellar sample with tight Teff and logg constraints to minimize systematic abundance trends and undertook a careful accounting of measurement errors. G23 assume that their data has negligible differential systematic scatter, and account only for the photon-noise scatter. Lacking abundances measurements from repeat observations of the same star, they use the error spectrum calculated in the PEPSI reduction pipeline to simulate 10 synthetic spectra for each star. They vary the flux at each wavelength by an amount drawn from a Gaussian with a standard deviation equivalent to the error. The new realizations of the spectra are used to recalculate the abundances and the standard deviation is adopted as the photon-noise. G23 find that the median photon-noise uncertainty is $\rm{\sigma_{phot,med}}<0.04$ dex. They also consider NLTE effects, but conclude that while the shape of the abundance pattern is influenced by NTLE effects, the abundance scatter is not.
G23 find intrinsic scatters ranging from $0.04-0.08$ dex in [X/Fe] for element abundances in our set (Mg, Si, Ti, Ni, Mn). This is greater than the photon-noise uncertainty by a factor of 2 for most elements. Overall, because the median photon-noise uncertainties are smaller than the scatter, small inaccuracies in the abundance uncertainty will have little impact on the intrinsic scatter. Under the assumption that these errors may be underestimated due to the assumption of negligible differential systematic scatter, the method provides robust upper limits on the intrinsic scatter of the element abundances for this population. If on the other hand, uncertainties on the data were underestimated by, for example, a factor of 2, intrinsic scatters would only increase by 0.01 dex for each element in our sample, well within the uncertainty on the measurement.
25 < S/N < 125 125 < S/N < 225 225 < S/N < 325 325 < S/N < 425 425 < S/N
[Al/Fe], [O/Fe] -2.1 < [Fe/H] < -1.55 $0.36\pm 0.10$ $0.23\pm 0.12$ $0.19\pm 0.09$ $0.15\pm 0.11$ $0.09\pm 0.07$
-1.55 < [Fe/H] < -1 $0.23\pm 0.03$ $0.08\pm 0.04$ $0.06\pm 0.06$ $0.04\pm 0.02$ $0.08\pm 0.04$
[Mg/Fe] -2.1 < [Fe/H] < -1.55 $0.46\pm 0.15$ $0.04\pm 0.03$ $0.05\pm 0.04$ $0.04\pm 0.03$ $0.04\pm 0.03$
-1.55 < [Fe/H] < -1 $0.24\pm 0.03$ $0.01\pm 0.01$ $0.01\pm 0.01$ $0.02\pm 0.02$ $0.04\pm 0.02$
[Si/Fe] -2.1 < [Fe/H] < -1.55 $0.43\pm 0.13$ $0.22\pm 0.08$ $0.24\pm 0.07$ $0.03\pm 0.03$ $0.02\pm 0.02$
-1.55 < [Fe/H] < -1 $0.26\pm 0.03$ $0.09\pm 0.02$ $0.04\pm 0.02$ $0.03\pm 0.02$ $0.08\pm 0.01$
[Ti/Fe] -2.1 < [Fe/H] < -1.55 $0.42\pm 0.22$ $0.19\pm 0.11$ $0.32\pm 0.11$ $0.31\pm 0.22$ $0.22\pm 0.05$
-1.55 < [Fe/H] < -1 $0.20\pm 0.03$ $0.13\pm 0.02$ $0.06\pm 0.02$ $0.05\pm 0.03$ $0.14\pm 0.06$
[Ni/Fe] -2.1 < [Fe/H] < -1.55 $0.17\pm 0.07$ $0.36\pm 0.13$ $0.16\pm 0.05$ $0.15\pm 0.06$ $0.05\pm 0.02$
-1.55 < [Fe/H] < -1 $0.18\pm 0.02$ $0.05\pm 0.01$ $0.02\pm 0.01$ $0.02\pm 0.01$ $0.03\pm 0.01$
[Mn/Fe] -2.1 < [Fe/H] < -1.55 $0.32\pm 0.11$ $0.36\pm 0.18$ $0.20\pm 0.07$ $0.25\pm 0.13$ $0.14\pm 0.04$
-1.55 < [Fe/H] < -1 $0.25\pm 0.03$ $0.11\pm 0.02$ $0.17\pm 0.02$ $0.08\pm 0.02$ $0.07\pm 0.01$
To be consistent with the calculated APOGEE intrinsic scatters, we use Equation 3 to calculate the intrinsic scatter of the G23 data for the 86 stars in their sample, showing the result in Figure 5. We report a range of intrinsic scatters in [X/Fe] of 0.04 -- 0.07 dex for elements that overlap with APOGEE (Mg, Si, Ti, Ni, Mn). The calculated intrinsic scatters for G23 are within (Mg, Si, Ti) or slightly lower (Ni, Mn) than the reported intrinsic scatters relative to the two- and three-parameter models reported in Table 4 of G23.
@@ 3.2.3 Calibration of APOGEE Uncertainties
Due to the robustness of intrinsic scatter from the G23 data, we take these measurements as the "ground truth" intrinsic scatter of halo field stars and calibrate the uncertainties on the APOGEE data such that the calculated intrinsic scatter of the APOGEE halo data matches that of G23. In practice, we calculate an additional term, $a$, for every abundance and add it in quadrature to the reported uncertainty on abundance measurements in APOGEE.
@@
Using equation 4 we fix the intrinsic scatter, $s$, to the calculated value from the G23 data and maximize the likelihood over $a$. We bin over the APOGEE halo data in [Fe/H] and S/N as we expect that the total uncertainties are primarily dependent on these quantities. As the morphology of the absorption features may also be influenced by the effective temperature, a better calibration of the errors may be obtained by binning with respect to $\rm{T_{eff}}$, however we are unable to do this robustly with our sample size. Due to the small number of high S/N stars at low metallicity, we divide our [Fe/H] range into two bins from $-2.1$ dex to $-1.55$ dex and $-1.55$ dex to $-1$ dex. We then bin the S/N in ranges of 100, from a S/N of 25 up to 425. The final bin contains all stars with S/N higher than 425 as there are very few stars in this range and S/N beyond this range are thought to be unphysically high due to persistence effects (Holtzman et al., 2018), though the data does still have very high S/N. We note that finer binning would result in more precise added error terms on the data, but again, we are limited by the APOGEE halo sample size. The median number of stars in each bin is listed in Table 1.
@@
Figure 6: Added error terms for all elemental abundances, binned by S/N and [Fe/H]. Within each elemental abundance, added error terms generally decrease with increasing S/N and [Fe/H].
Figure 6 shows the added error terms for all abundances explored with the APOGEE data. There is a clear trend of decreasing added error terms with higher S/N and [Fe/H]. This is expected as higher S/N and higher metallicity data have stronger metal lines. Furthermore, we note that the size of the added error terms at low S/N are on the order of the uncalibrated intrinsic scatter. This implies the intrinsic scatters are very sensitive to the estimated uncertainty which is a limitation of the data quality.
As the intrinsic scatters reported by G23 are potentially upper limits themselves, the intrinsic scatters measured in this study, calibrated by the added error term, are (while measurements) also potentially upper bounds as they are limited by the accuracy on the uncertainties of the G23 data. Having constructed the G23 halo and APOGEE samples to be drawn from nearly the same underlying population, this is a robust data-driven method to obtain accurate uncertainties that we can propagate for our analysis of the dSph systems and provides a benchmark for comparison for further studies.
@@ 4.1 Milky Way Stellar Halo
To validate our calibrated uncertainty estimates, we first compare the calibrated intrinsic scatter of the APOGEE halo sample to the G23 halo sample as shown in Figure 7. Calibrated intrinsic scatters for all abundances decrease by at least 0.02 dex compared to the uncalibrated intrinsic scatters. For those abundances overlapping with the G23 data, intrinsic scatters for the calibrated APOGEE data are approximately the same as the intrinsic scatters derived from the G23 data, as expected. Specifically, we report intrinsic scatters of 0.07 dex for [Si/Fe], 0.04 dex for [Ti/Fe], 0.04 dex for [Ni/Fe], and 0.06 for [Mn/Fe] which are all within 0.01 dex of G23. We report an intrinsic scatter of 0.05 for [Mg/Fe], notably nearly 0.02 dex lower than the G23 value. This suggests that the added error term may be too large for some data and that the data may need to be more finely binned to get an accurate estimate of the added error for each star, which would require a larger sample of low-metallicity stars. In general, the differences compared to
G23 are likely due to the coarse binning for the added error term. In addition, since the maximization of the joint probability function accounts for the individual errors on stars, and the added error terms are not calibrated for individual stars; this likely introduces some small differences between the intrinsic scatter in the G23 data and the calibrated APOGEE data. Nevertheless, these differences are slight and the comparison in Figure 7 validates our procedure.
@@ 4.2 Milky Way dSphs
With the robustness of our calibration confirmed, we now turn to an analysis of the dSphs. Figure 8 shows the comparison between the calibrated and uncalibrated average dSph intrinsic scatter, calibrated intrinsic scatters for individual dSphs in APOGEE, and the average intrinsic scatters from the K18 sample. The calibrated average intrinsic scatter of dSphs is reduced by up to 0.10 dex over the uncalibrated data, making the average scatter of dSphs in the APOGEE data comparable to the average scatter of dSphs in K18.
[…]
@@ 4.3 Milky Way Globular Clusters
Figure 9 shows the average intrinsic scatters of 1604 stars across 19 GCs (median 79 stars per GC) that were targeted in APOGEE DR17 (see the Appendix for a list of GCs). We separately analyze Omega Centauri and M54, as they are suspected cores of disrupted dwarf galaxies (Neumayer et al., 2020). The GCs have a median S/N of 157, and [Fe/H] of -1.1. The average intrinsic scatter is 0.11 dex for [Al/Fe] and 0.07 dex for [O/Fe]; higher than the $\alpha$-elements and Fe-peak elements, each of which have an average scatter of 0.04 dex. This is consistent with the multiple stellar populations mentioned previously. We also note that our results are consistent with abundance dispersions previously measured in GCs (e.g. Yong et al., 2013).
@@ 5.1 Summary
In this work we introduce a methodology for calibrating abundance uncertainties derived from mid-resolution surveys by leveraging smaller, high-fidelity, high-resolution data sets. Data-driven calibration or label-transfer efforts have previously been employed to garner greater value from low S/N data (e.g. Casey et al., 2017; Ho et al., 2017; Rice & Brewer, 2020; Xiang et al., 2021, 2022). However, these typically rely on stars in common between the surveys. Our approach does not rely on this, but instead utilizes population statistics. We apply our method to calibrate the measurement errors and calculate the intrinsic abundance scatter of metal-poor stars for elements Al, O, Mg, Si, Ti, Ni, and Mn in dSphs, GCs, and the Milky Way stellar halo. Our calculation relies on the calibration of the reported APOGEE abundance uncertainties and is specifically for APOGEE DR17 metal-poor red giant stars ($-2.1\leq$[Fe/H]$\leq-1$) and in the temperature range 4000KT $4000\,\rm{K}\,\leq T\leq\,5000\,\rm{K}$.
@@
Making measurements of the intrinsic abundance scatter of dSphs, GCs, and the halo with the APOGEE data, particularly for metal-poor stars, is important for providing constraints for simulations and linking to physical drivers. Doing this requires using data at lower metallicities and S/N than the majority of APOGEE spectra, in a regime where there is little previous work. As the magnitude of the uncertainties impact the measurement of intrinsic scatter, it is vital to test the accuracy of the APOGEE reported uncertainty. In doing so, we found them to be underestimated and using the method introduced in this paper, we derive the complete uncertainties on the APOGEE data in this regime.
@@
To do this, we use a high fidelity reference set of halo stars (G23) to provide ground truth information about the intrinsic scatter of element abundances in the stellar halo. We construct a comparable sample of field halo stars from the APOGEE data using the same distance and metallicity criteria as G23. Our primary assumption in this work is that the underlying population of stars in the reference set (G23) is comparable to the halo population in the calibration set we use from APOGEE. Comparing the intrinsic scatters between the two datasets, we find that the intrinsic scatters of the uncalibrated APOGEE halo data are substantially higher than G23, indicating the APOGEE uncertainties are underestimated. This is unsurprising as the reference objects used in APOGEE for providing empirical uncertainty estimates are sparse in the metal-poor and low S/N regime (S/N $\sim$ 50) (Jönsson et al., 2020). We note that the G23 uncertainties themselves are very small so their intrinsic scatters, which we take as ground truth, are minimally impacted by any inaccuracy in these. However, strictly, any underestimation in the G23 uncertainties would result in lower intrinsic scatter, and so these data set an upper limit on the ground truth intrinsic scatter and therefore a lower limit on the complete uncertainties for APOGEE.
To calibrate the APOGEE uncertainties, we determine the uncertainty on the APOGEE halo data required to match the intrinsic abundance scatter of the halo population from G23. We subsequently calculate an added error term for every abundance as a function of [Fe/H] binned across a range of -2.1 to -1 and S/N binned across a range of 25 to 1420, which we can apply to stars in this parameter space.
After calibrating the uncertainties for each of the stars in the dSphs, we find that on average, dSphs have comparatively low intrinsic scatter in the light and $\alpha$-elements (on average, 0.04 dex) and higher intrinsic scatter in the Fe-peak elements (on average, 0.07 dex). We also undertake a similar analysis for GCs, but show that in contrast to the dSphs, GCs have high intrinsic scatter in the light elements (Al and O with 0.11 and 0.07 dex respectively), and no clear pattern in intrinsic scatter among the $\alpha$- and Fe-peak elements. This may suggest that for elements ejected through higher-energy processes, mixing is most efficient on small scales.
5.2 Comparison of dSphs, MW Halo, and GCs
Comparing the average intrinsic abundance scatters of dSphs, GCs, and the stellar halo enables us to set constraints on the origin of scatter and the formation histories of different structures. Figure 10 shows that the intrinsic scatter of the halo is more similar to the average scatter of GCs than it is to the average scatter of dSphs. However, this is not to necessarily say that the halo and GCs are formed of the same populations of stars, rather, that we distinctly see the signatures of multiple populations in both. This may be an indicator of the existence of stars from destroyed GCs in the halo, a theory supported by the high abundance dispersions in Al, O, and Si over Fe found in the in situ halo (Belokurov & Kravtsov, 2022) which have been linked to an increase contribution from GCs (Belokurov & Kravtsov, 2023). This may also be a signature of the combination of dwarfs with multiple populations of different element abundance means and scatters, as we expect for the composition of the halo, which would contribute to a larger scatter.
Additionally, that the classical dSphs have a different chemical composition than the halo has been shown both in observational (Unavane et al., 1996; Venn et al., 2004) and simulated data (Robertson et al., 2005; Font et al., 2006; Johnston et al., 2008; Cunningham et al., 2022; Horta et al., 2023b). Much of the difference is attributed to the short star formation histories of accreted dwarfs compared to longer-lived classical dSphs. The higher scatter seen in the halo could be interpreted as a result of early quenching of star formation in dwarfs destroyed by the halo, where metals do not have time to mix and end up in stars before star formation ceases. In comparison, the classical dSphs, such as those in our sample have more extended star formation histories which allow for more complete mixing, resulting in lower intrinsic scatters.
5.3 Future Prospects for Observations and Simulations
The ability to account for the complete uncertainties in the metal-poor regime of survey data is critical in order to fully exploit these data. Large-scale surveys also offer unique coverage of wider areas of dSphs than smaller studies, which typically focus on the core. Broader area coverage in general enables studies of element abundance scatter across the Galaxy, as well as spatially dependent scatter. Currently, available survey data of dSphs stars are generally low S/N ($<70$ median in APOGEE), are limited to the brightest stars, and element abundance measurement uncertainties are nearly as high as the total abundance scatter. In order to make high-fidelity measurements of the intrinsic scatter of element abundances of low-metallicity dSph systems, reasonable sample sizes of high resolution (e.g. R $>$ 50,000), high S/N spectra (e.g. $>$ 100) are required.
However, our efforts to calibrate large medium-resolution survey data using small high resolution, high S/N studies makes the most use of the full parameter space of the available data, as well as provides a better understanding of the unaccounted uncertainties in the metal-poor regime in APOGEE, which is relevant to an ensemble of studies. Our results subsequently serve as a baseline for the next generation of data. High-resolution, high S/N spectra can be obtained with upcoming 30-meter class telescope programs such as the Giant Magellan Telescope (GMT), European Extremely Large Telescope (E-ELT), and Thirty Meter Telescope (TMT). Additionally, 30-meter class telescopes are expected to reach up to the 25th magnitude in the V-band, or 2 magnitudes deeper than existing observations of resolved stars in dSphs (Ji et al., 2019). The intrinsic scatter in element abundances as we showcase here can serve as an important test and comparison for future ELT studies where measurement uncertainties will undoubtedly be smaller, and will serve as another source of calibration data for the uncertainty estimates used in this work.